# Multispectral airborne laser scanning for tree species classification: a benchmark of machine learning and deep learning algorithms

### v1

@@ 3.2.1 Point-based (3D) deep learning methods
SLU-YOLOv8-DL-2D, HeliALS For each tree, points from VUX and mini-VUX were merged and the point cloud was projected to four 2D RGB images corresponding to different rotations around the z-axis (0, 45, 90, or 135). Green channel was not used. $4\times\mathrm{2D}$ RGB images ($160\times 320$ pixels) per tree. For each pixel, the RGB values correspond to a mean mini-VUX intensity (NIR), mean VUX intensity (SWIR), and a combination of the two, respectively. YOLOv8 (Jocher et al., 2023) pretrained on the ImageNet dataset (Deng et al., 2009) was finetuned over 15 training epochs. The RGB images for rotation angles 0, 45, and 90 were used in the training set, while the RGB images for the angle 135 formed the validation set. Trained YOLOv8 model predicted the species for each of the four RGB images generated from a single tree, and the results were aggregated to obtain the final prediction.
@@
DetailView-DL-2D, HeliALS Removal of noise and ground points. Voxelization with a side length of 5 cm to obtain a more balanced point distribution. Point clouds of the three channels merged. Intensity and reflectance ignored. $(4+1+1+1)\times\mathrm{2D}$ depth images ($256\times 256$ pixels) per tree. Intensity information was not used. Out of the seven depth images, four were from the sides, one from the top, one from the bottom, and one representing a close-up between $z=1.0$ m and $z=1.5$ m. DetailView neural net model (Puliti et al., 2025; Frey and Schindler, 2024) based on DenseNet-201-instances (Huang et al., 2017) trained from scratch over 24 epochs using 90-10 train-validation split. The batch size was set twice as large and the learning rate a factor of 10 smaller compared to Puliti et al. (2025). Augmentations included point cloud random subsampling and rotations around the z-axis. Predicted class probabilities averaged across 50 augmentations of each tree.
@@
SA-Detailview-DL-2D, HeliALS Point clouds of mini-VUX (NIR) and VUX (SWIR) merged. Removal of ground points. The green channel was ignored. Same as DetailView-DL-2D. DetailView neural net model (Puliti et al., 2025; Frey and Schindler, 2024) based on DenseNet-201-instances (Huang et al., 2017). The model had been previously trained on the FOR-species dataset with 33 different species and 98-2 training-validation split using same augmentations as DetailView-DL-2D. The model was NOT re-trained using the provided training set. Predicted class probabilities averaged across 50 augmentations of the tree. If predicted tree species was not present in the current dataset, it was reclassified to a species in the dataset based on the largest relative accuracy in the confusion matrix.
@@
NTNU-ConvNeXt-T-DL-2D, Optech Titan Point clouds of the three scanners merged. For each tree, the point cloud was converted to 6 depth images corresponding to 4 different rotations around the z-axis and 2 projections along the negative and positive z-axis following Goyal et al. (2021). Furthermore, a front-view depth image was obtained between $z=1.0$ m and $z=2.5$ m. $(6+1)\times\mathrm{2D}$ depth images ($128\times 128$ pixels that was rescaled before input layer to $256\times 256$ pixels) per tree. Furthermore, 20 additional global features were computed for each tree based on the tree height and intensities of the 3 channels following Yu et al. (2017). Single-channel intensity features included, e.g., minimum, maximum, and moments. Neural net with 3 input branches corresponding to the 6 depth images from the sides, the front-view depth image, and the global features. ConvNeXt-T (Liu et al., 2022b) provided the backbone for the two image branches, and a multi-layer perceptron block for the global features. Training-validation split of 80-20 with data augmentations consisting of random rotations around the z-axis, translations and scaling. Prediction of the trained model.
@@ 3.2.2 Profile-based (2D) deep learning methods
Our comparison of 2D deep learning approaches includes four methods: SLU-YOLOv8-DL-2D, DetailView-DL-2D, SA-Detailview-DL-2D, and NTNU-ConvNeXt-T-DL-2D. Importantly, SA-Detailview-DL-2D represents a sensor agnostic (SA) deep learning model that has been previously trained on the FOR-species20K dataset (Puliti et al., 2024) without any re-training on the provided training dataset. All the other models, including Detailview-DL-2D, have been trained on the training set provided as a part of the benchmarking competition. When it comes to the used radiometric features, NTNU-ConvNeXt-T-DL-2D utilized radiometric information from all of the three channels. SLU-YOLOv8-DL-2D took into account both geometry and radiometric information from VUX and mini-VUX scanners, while ignoring the data of the green channel. In contrast, DetailView utilized only the geometry of the multispectral point cloud without any radiometric information.
The studied 2D deep learning methods are based on various convolutional neural net backbones, including YOLOv8 (Jocher et al., 2023) for SLU-YOLOv8-DL-2D, DenseNet-201 (Huang et al., 2017) for DetailView, and ConvNeXt-T (Liu et al., 2022b) for NTNU-ConvNeXt-T-DL-2D. DetailView and NTNU-ConvNeXt-T-DL-2D convert the segment-wise point clouds into multiple depth image projections from different angles. On the other hand, SLU-YOLOv8-DL-2D projects the point clouds into multiple RGB images, where the RGB values of each pixel are based on the mean intensities of VUX, mini-VUX and their combination. As their input, all the models take between four and seven 2D projections from different viewing angles. The resolutions of the projections range from $128\times 128$ pixels for NTNU-ConvNeXt-T-DL-2D to $256\times 256$ pixels for DetailView and $160\times 320$ pixels for SLU-YOLOv8-DL-2D. NTNU-ConvNeXt-T-DL-2D rescales the point cloud projection to $256\times 256$ pixels before the input layer of the model. All of the methods use side-view projections corresponding to different rotations around the z-axis. Furthermore, DetailView utilizes a top-view projection, a bottom-view projection, and a projection to a close-up of the trunk. In addition to the side-, top- and bottom-view projections, NTNU-ConvNeXt-T-DL-2D includes a front-view depth image and 20 additional global features mostly based on the intensities of the three channels. While most of the methods utilize orthographic projections, NTNU-ConvNeXt-T-DL-2D follows the approach by Goyal et al. (2021) and performs the projection using a pinhole camera model (perspective projection).
@@
IBL-BalancedRF-ML, Optech Titan Segments with under- or over-segmentation errors were identified and removed, which resulted in 660 training trees. Following Kamińska et al. (2021) and Lisiewicz et al. (2025), features were computed, including geometric features based on the merged point cloud, and intensity features based on individual channels and their combinations. For most features, calculations were performed for points situated above half the height of each segment. A feature vector of 30 elements after dimensionality reduction. Variables based on combinations of different channels had the highest importance. Random forest model was trained with a balanced RF approach (Chen et al., 2004), up-sampling minority species. To ensure stable results and mitigate overfitting, 5-fold cross-validation was repeated 20 times.
UEF-RF/LGBM/SVM-ML, Optech Titan Following Holmgren and Persson (2004), Ørka et al. (2009), Ørka et al. (2012), and Hovi et al. (2016), intensity features were computed, including single-channel moments as a function of height for each echo type, and two multi-channel features known as the normalized difference water index and conifer index (Trier et al., 2018). The only geometric feature was the 95th percentile of the $z$ coordinate. A feature vector of 143 elements after dimensionality reduction and 259 elements prior to dimensionality reduction. In the dimensionality reduction, correlation coefficient was evaluated for each pairs of features, and one of the features in the pair was removed if the correlation coefficient exceeded 0.9. Three ML methods were tested for the classification, including random forest, light gradient boosting machine (LGBM), and support vector machine (SVM) using radial basis function (RBF) kernel. 80-20 train-test split used. Grid search and cross validation were utilized to optimize hyperparameters. No augmentation or species imbalance mitigation.
@@
FGI-RF-ML, HeliALS and Optech Titan 92 features computed for each tree before dimensionality reduction. Following Lehtomäki et al. (2016), Yu et al. (2017), and Hakula et al. (2023), features included geometric features using the channel with most points, intensity statistics including, e.g., normalized intensity histograms for each channel, proportions of laser pulse return types (e.g. first of many), and histogram of local descriptors. A feature vector of 92 elements before dimensionality reduction and 20–40 elements after dimensionality reduction. Random forest classifier trained based on out-of-bag prediction. During training, feature selection and model optimization were performed using a grid search based on the out-of-bag prediction. No augmentation or species imbalance mitigation.
@@
LUKE-MultiRF-ML, Optech Titan 36 features computed for each tree, including, e.g., moments and quantiles of the $z$-coordinate and the radial distance from the center of the tree, average intensities of the channels, average intensities of the first and other echoes for each channel, and features combining intensities of two channels. A feature vector of 36 elements. 10 random forest models trained using 5-fold cross validation, including a default model classifying trees into the nine species and nine single-species models. Synthetic minority over-sampling technique (Chawla et al., 2002) used to balance the training set with the oversampling parameter set to 10 for pine, spruce, and birch, and to 15 for the rest. If exactly one of the single-species classifiers predicted the tree to belong to its species at inference time, the tree was assigned to this species. Otherwise, the species of the tree was assigned based on the default model.
@@
UPV-GB-ML, Optech Titan Point clouds were filtered to remove ground points and outliers. In total, 384 geometric features and 48 single-channel intensity features (16 for each channel) were computed. Single-channel intensity features included, e.g., min, mean, standard deviation and different percentiles. Similar statistical metrics were evaluated for geometrical neighborhood features, all included in the Class3Dp software (Carbonell-Rivera et al., 2024), such as, number of returns, eigenvalues of the covariance matrix, and features based on West et al. (2004), Rusu (2010), and Demantké et al. (2012). A feature vector of 18 elements after dimensionality reduction. The dimensionality reduction was based on ranking the features based on their permutation importance and then evaluating the learning curve. The learning curve plateaued after 18 features. Gradient boosting (GB) machine used for classification. The training was performed using 5-fold cross validation. No augmentation or species imbalance mitigation.
@@
Aalto-RF-ML, Optech Titan 34 features computed for each tree, including e.g., intensity statistics for each channel based on points within the top 3 m of the tree, crown diameter within and below the top 3 m of the tree, the apex angle of the tree, and features combining intensities of two channels following Hakula et al. (2023). A feature vector of 34 elements. Random forest classifier trained using 80-20 training validation split. To mitigate class imbalance, new feature vectors were simulated for each species with fewer than 125 examples using the class-wise mean and standard deviation of the feature vector (Rönnholm et al., 2022). After augmentation, each class had at least 125 training examples.
@@
To generate more training examples, DetailView utilizes point cloud random subsampling and rotations around the z-axis as augmentations. SLU-YOLOv8-DL-2D utilizes four rotational augmentations around the z-axis per tree, while NTNU-ConvNeXt-T-DL-2D applies random translation and scaling in addition to random rotations around the z-axis. At inference time, SLU-YOLOv8-DL-2D and DetailView use an ensemble voting strategy to obtain the final prediction, whereas NTNU-ConvNeXt-T-DL-2D directly uses the output of the model for prediction. DetailView performs the prediction for 50 different augmentations of the tree increasing the required amount of computation.
@@ 3.2.3 Machine learning methods
Our comparison includes the following methods: IBL-BalancedRF-ML, UEF-RF/LGBM/SVM-ML, IDEAS-RF-ML, FGI-RF-ML, LUKE-MultiRF-ML, UPV-GB-ML, and Aalto-RF-ML. Most of the studied methods, such as IBL-BalancedRF-ML, UEF-RF-ML, IDEAS-RF-ML, FGI-RF-ML, LUKE-MultiRF-ML, and Aalto-RF-ML, utilize a random forest classifier or its variant for the species prediction. IBL-BalancedRF-ML applies a balanced RF procedure (Chen et al., 2004) to up-sample the number of training examples for minority species, whereas LUKE-MultiRF-ML utilizes multiple RF models for the species prediction. In addition to the RF classifiers, our comparison includes a support vector machine (SVM) used by UEF-SVM-ML, and variants of gradient boosting (GB) techniques used by UEF-LGBM-ML and UPV-GB-ML.
@@ 3.3 Accuracy analysis
[…]
@@
Moreover, we used the classification error as a performance metric when appropriate: the overall classification error refers to the complement of the overall accuracy, whereas the macro-average classification error is the complement of the macro-average accuracy.

### v2

@@ 3.2.1 Point-based (3D) deep learning methods
SLU-YOLOv8-DL-2D, HeliALS For each segment, points from VUX and mini-VUX were merged and the point cloud was projected to four 2D RGB images corresponding to different rotations around the z-axis (0, 45, 90, or 135). Green channel was not used. $4\times\mathrm{2D}$ RGB images ($160\times 320$ pixels) per segment. For each pixel, the RGB values correspond to a mean mini-VUX intensity (NIR), mean VUX intensity (SWIR), and a combination of the two, respectively. YOLOv8 (Yolov82023) pretrained on the ImageNet dataset (deng2009imagenet) was finetuned over 15 training epochs. The RGB images for rotation angles 0, 45, and 90 were used in the training set, while the RGB images for the angle 135 formed the validation set. Trained YOLOv8 model predicted the species for each of the four RGB images generated from a single segment, and the results were aggregated to obtain the final prediction.
@@
DetailView-DL-2D, HeliALS Removal of noise and ground points. Voxelization with a side length of 5 cm to obtain a more balanced point distribution. Point clouds of the three channels merged. Intensity and reflectance ignored. $(4+1+1+1)\times\mathrm{2D}$ depth images ($256\times 256$ pixels) per segment. Intensity information was not used. Out of the seven depth images, four were from the sides, one from the top, one from the bottom, and one representing a close-up between $z=1.0$ m and $z=1.5$ m. DetailView neural net model (puliti2025benchmarking; Detailview2024) based on DenseNet-201-instances (huang2017densely) trained from scratch over 24 epochs using 90-10 train-validation split. The batch size was set twice as large and the learning rate a factor of 10 smaller compared to puliti2025benchmarking. Augmentations included point cloud random subsampling and rotations around the z-axis. Predicted class probabilities averaged across 50 augmentations of each segment.
@@
SA-Detailview-DL-2D, HeliALS Point clouds of mini-VUX (NIR) and VUX (SWIR) merged. Removal of ground points. The green channel was ignored. Same as DetailView-DL-2D. DetailView neural net model (puliti2025benchmarking; Detailview2024) based on DenseNet-201-instances (huang2017densely). The model had been previously trained on the FOR-species dataset with 33 different species and 98-2 training-validation split using same augmentations as DetailView-DL-2D. The model was NOT re-trained using the provided training set. Predicted class probabilities averaged across 50 augmentations of the segment. If predicted tree species was not present in the current dataset, it was reclassified to a species in the dataset based on the largest relative accuracy in the confusion matrix.
@@
NTNU-ConvNeXt-T-DL-2D, Optech Titan Point clouds of the three scanners merged. For each segment, the point cloud was converted to 6 depth images corresponding to 4 different rotations around the z-axis and 2 projections along the negative and positive z-axis following goyal2021revisiting. Furthermore, a front-view depth image was obtained between $z=1.0$ m and $z=2.5$ m. $(6+1)\times\mathrm{2D}$ depth images ($128\times 128$ pixels that was rescaled before input layer to $256\times 256$ pixels) per segment. Furthermore, 20 additional global features were computed for each segment based on the tree height and intensities of the 3 channels following yu2017single. Single-channel intensity features included, e.g., minimum, maximum, and moments. Neural net with 3 input branches corresponding to the 6 depth images from the sides, the front-view depth image, and the global features. ConvNeXt-T (liu2022convnet) provided the backbone for the two image branches, and a multi-layer perceptron block for the global features. Training-validation split of 80-20 with data augmentations consisting of random rotations around the z-axis, translations and scaling. Prediction of the trained model.
@@ 3.2.2 Profile-based (2D) deep learning methods
Our comparison of 2D deep learning approaches includes four methods: SLU-YOLOv8-DL-2D, DetailView-DL-2D, SA-Detailview-DL-2D, and NTNU-ConvNeXt-T-DL-2D. We note that SA-Detailview-DL-2D represents a sensor agnostic (SA) deep learning model that has been previously trained on the FOR-species20K dataset (puliti_2024_13255198) without any re-training on the provided training dataset. All the other models, including Detailview-DL-2D, have been trained on the training set provided as a part of the benchmarking competition. When it comes to the used spectral features, NTNU-ConvNeXt-T-DL-2D utilized radiometric information from all of the three channels. SLU-YOLOv8-DL-2D took into account both geometry and spectral information from VUX and mini-VUX scanners, while ignoring the data of the green channel. In contrast, DetailView utilized only the geometry of the multispectral point cloud without any radiometric information.
The studied 2D deep learning methods are based on various convolutional neural net backbones, including YOLOv8 (Yolov82023) for SLU-YOLOv8-DL-2D, DenseNet-201 (huang2017densely) for DetailView, and ConvNeXt-T (liu2022convnet) for NTNU-ConvNeXt-T-DL-2D. DetailView and NTNU-ConvNeXt-T-DL-2D convert the segment-wise point clouds into multiple depth image projections from different angles. On the other hand, SLU-YOLOv8-DL-2D projects the point clouds into multiple RGB images, where the RGB values of each pixel are based on the mean intensities of VUX, mini-VUX and their combination. As their input, all the models take between four and seven 2D projections from different viewing angles. The resolutions of the projections range from $128\times 128$ pixels for NTNU-ConvNeXt-T-DL-2D to $256\times 256$ pixels for DetailView and $160\times 320$ pixels for SLU-YOLOv8-DL-2D. NTNU-ConvNeXt-T-DL-2D rescales the point cloud projection to $256\times 256$ pixels before the input layer of the model. All of the methods use side-view projections corresponding to different rotations around the z-axis. Furthermore, DetailView utilizes a top-view projection, a bottom-view projection, and a projection to a close-up of the trunk. In addition to the side-, top- and bottom-view projections, NTNU-ConvNeXt-T-DL-2D includes a front-view depth image and 20 additional global features mostly based on the intensities of the three channels. While most of the methods utilize orthographic projections, NTNU-ConvNeXt-T-DL-2D follows the approach by goyal2021revisiting and performs the projection using a pinhole camera model (perspective projection).
@@
IBL-BalancedRF-ML, Optech Titan Segments with under- or over-segmentation errors were identified and removed, which resulted in 660 training segments. Following kaminska2021single and lisiewicz2025comprehensive, features were computed, including geometric features based on the merged point cloud, and intensity features based on individual channels and their combinations. For most features, calculations were performed for points situated above half the height of each segment. A feature vector of 30 elements after dimensionality reduction. Variables based on combinations of different channels had the highest importance. Random forest model was trained with a balanced RF approach (chen2004using), up-sampling minority species. To ensure stable results and mitigate overfitting, 5-fold cross-validation was repeated 20 times.
UEF-RF/LGBM/SVM-ML, Optech Titan Following holmgren2004identifying, orka2009classifying, orka2012simultaneously, and hovi2016lidar, intensity features were computed, including single-channel moments as a function of height for each echo type, and two multi-channel features known as the normalized difference water index and conifer index (trier2018tree). The only geometric feature was the 95th percentile of the $z$ coordinate. A feature vector of 143 elements after dimensionality reduction and 259 elements prior to dimensionality reduction. In the dimensionality reduction, correlation coefficient was evaluated for each pairs of features, and one of the features in the pair was removed if the correlation coefficient exceeded 0.9. Three ML methods were tested for the classification, including random forest, light gradient boosting machine (LGBM), and support vector machine (SVM) using radial basis function (RBF) kernel. 80-20 train-test split used. Grid search and cross validation were utilized to optimize hyperparameters. No augmentation or species imbalance mitigation.
@@
FGI-RF-ML, HeliALS and Optech Titan 92 features computed for each segment before dimensionality reduction. Following lehtomaki2015object, yu2017single, and hakula2023individual, features included geometric features using the channel with most points, intensity statistics including, e.g., normalized intensity histograms for each channel, proportions of laser pulse return types (e.g. first of many), and histogram of local descriptors. A feature vector of 92 elements before dimensionality reduction and 20–40 elements after dimensionality reduction. Random forest classifier trained based on out-of-bag prediction. During training, feature selection and model optimization were performed using a grid search based on the out-of-bag prediction. No augmentation or species imbalance mitigation.
@@
LUKE-MultiRF-ML, Optech Titan 36 features computed for each segment, including, e.g., moments and quantiles of the $z$-coordinate and the radial distance from the center of the segment, average intensities of the channels, average intensities of the first and other echoes for each channel, and features combining intensities of two channels. A feature vector of 36 elements. 10 random forest models trained using 5-fold cross validation, including a default model classifying segments into the nine species and nine single-species models. Synthetic minority over-sampling technique (chawla2002smote) used to balance the training set with the oversampling parameter set to 10 for pine, spruce, and birch, and to 15 for the rest. If exactly one of the single-species classifiers predicted the segment to belong to its species at inference time, the segment was assigned to this species. Otherwise, the species of the segment was assigned based on the default model.
@@
UPV-GB-ML, Optech Titan Point clouds were filtered to remove ground points and outliers. In total, 384 geometric features and 48 single-channel intensity features (16 for each channel) were computed. Single-channel intensity features included, e.g., min, mean, standard deviation and different percentiles. Similar statistical metrics were evaluated for geometrical neighborhood features, all included in the Class3Dp software (carbonell2024class3dp), such as, number of returns, eigenvalues of the covariance matrix, and features based on west2004context, rusu2010semantic, and demantke2012streamed. A feature vector of 18 elements after dimensionality reduction. The dimensionality reduction was based on ranking the features based on their permutation importance and then evaluating the learning curve. The learning curve plateaued after 18 features. Gradient boosting (GB) machine used for classification. The training was performed using 5-fold cross validation. No augmentation or species imbalance mitigation.
@@
Aalto-RF-ML, Optech Titan 34 features computed for each segment, including e.g., intensity statistics for each channel based on points within the top 3 m of the segment, crown diameter within and below the top 3 m of the segment, the apex angle of the tree, and features combining intensities of two channels following hakula2023individual. A feature vector of 34 elements. Random forest classifier trained using 80-20 training validation split. To mitigate class imbalance, new feature vectors were simulated for each species with fewer than 125 examples using the class-wise mean and standard deviation of the feature vector (ronnholm2022utilising). After augmentation, each class had at least 125 training examples.
@@
To generate more training examples, DetailView utilizes point cloud random subsampling and rotations around the z-axis as augmentations. SLU-YOLOv8-DL-2D utilizes four rotational augmentations around the z-axis per segment, while NTNU-ConvNeXt-T-DL-2D applies random translation and scaling in addition to random rotations around the z-axis. At inference time, SLU-YOLOv8-DL-2D and DetailView use an ensemble voting strategy to obtain the final prediction, whereas NTNU-ConvNeXt-T-DL-2D directly uses the output of the model for prediction. DetailView performs the prediction for 50 different augmentations of the segment increasing the required amount of computation.
@@ 3.2.3 Machine learning methods
Our comparison includes the following methods: IBL-BalancedRF-ML, UEF-RF/LGBM/SVM-ML, IDEAS-RF-ML, FGI-RF-ML, LUKE-MultiRF-ML, UPV-GB-ML, and Aalto-RF-ML. Most of the studied methods, such as IBL-BalancedRF-ML, UEF-RF-ML, IDEAS-RF-ML, FGI-RF-ML, LUKE-MultiRF-ML, and Aalto-RF-ML, utilize a random forest classifier or its variant for the species prediction. IBL-BalancedRF-ML applies a balanced RF procedure (chen2004using) to up-sample the number of training examples for minority species, whereas LUKE-MultiRF-ML utilizes multiple RF models for the species prediction. In addition to the RF classifiers, our comparison includes a support vector machine (SVM) used by UEF-SVM-ML, and variants of gradient boosting (GB) techniques used by UEF-LGBM-ML and UPV-GB-ML.
@@ 3.3 Accuracy analysis
$\text{Macro-average accuracy}=\frac{1}{|S|}\sum_{s\in S}\frac{\text{TP}_{s}}{\text{TP}_{s}+\text{FN}_{s}}$ (5)
@@
Moreover, we used the classification error as a performance metric when appropriate: the overall classification error refers to the complement of the overall accuracy, i.e.,
$\varepsilon_{\mathrm{OA}}=1-\mathrm{OA},$ (6)
where $\varepsilon_{\mathrm{OA}}$ denotes the overall classification error. Similarly, the macro-average classification error is the complement of the macro-average accuracy.
We further estimated the uncertainty of overall accuracy and macro-average accuracy for each of the classification methods using a bootstrapping approach. Based on the test set of our dataset, we generated 2000 bootstrapping datasets of 5261 tree segments using random sampling with replacement. For each classification method, we recalculated the accuracy metrics for each bootstrapping dataset, resulting in a bootstrapping distribution of 2000 accuracy values. We derived a 95% confidence interval (CI) by taking the 2.5th and 97.5th percentiles of the distribution. The estimated confidence intervals reflect the uncertainty arising from the finite test set, but do not include, e.g., randomness in the model training.
3.4 Scaling analysis of classification error
In addition to benchmarking the classification methods on the fixed training and test set introduced in Sec. 2.4, we also studied the scaling of the classification error as a function of training set size and point density. We conducted these experiments using the deep learning and machine learning models achieving the highest overall accuracy on the HeliALS dataset in the benchmark contest. In the scaling studies, we excluded 88 segments from the original training set of the benchmarking competition due to non-optimal profile images, sparse point clouds, or building intersection within the segment boundaries. In addition, the species label of 18 segments from the original training set was corrected due to errors detected during the on-site verification. Compared to the test set of the benchmark contest, the dataset for scaling studies included 19 additional segments. In total, the dataset for scaling law studies consisted of 6257 tree segments. Below, this dataset is called the scaling analysis dataset. Importantly, these modifications to the dataset did not significantly affect the classification results. For example, FGI-PointTransformer-DL-3D obtained an overall accuracy of 87.9% when trained on the original training set of 1065 segments in the benchmark competition. The same model obtained an overall accuracy of 88.5% when trained on 1000 segments from the scaling analysis dataset.
In the scaling studies, we estimated the overall and macro-average classification errors of the selected classification models using 5-fold cross-validation. The folds were selected using stratified random sampling to ensure that each fold follows the same species distribution. For the scaling experiments with varying training set size, we utilized 5-fold cross-validation on subsets of the scaling analysis dataset, thus effectively creating training sets ranging from approximately 250 segments to approximately 5000 segments. Importantly, each fold in the subsets had a similar species distribution as the original dataset as a result of another round of stratified random sub-sampling from the folds previously generated from the full dataset.
For the scaling experiments as a function of point density, we randomly sub-sampled the HeliALS tree segments to a desired point density. We varied the point density from $10^{-1}$ $\mathrm{pts}/\mathrm{m}^{2}$ representing an extremely sparse point cloud to around 1000 $\mathrm{pts}/\mathrm{m}^{2}$ representing a dense ALS point cloud. We evaluated the classification performance using 5-fold cross-validation on the entire sub-sampled scaling analysis dataset.